OpenStudy (r.hamza17):

At 10% C.I. p.a., a sum of money accumulate to Rs. 9625 in 5 years. The sum initially invested is?

OpenStudy (trojanpoem):

Let the initial sum d: The sum at any year (y) with C.I.P.A (c) and final accumulate (s) is given by: \[d + \frac{ c }{ 100 } d * y = S\] \[d + \frac{ 10 }{ 100 } d * 5= 9625\] \[1.5d = 9625\]

OpenStudy (trojanpoem):

If you started with d money and you get each year 10 percent of that amount so finally you will have your inital investment + amount gained each year we gain 10 percent = 0.1d (10/100) d + 0.1d * (years) = total now.

OpenStudy (r.hamza17):

The answer I am getting through this method is not right.. The options given are: a) 5976.37 b) 5970 c) 5975 d) 5370.96

OpenStudy (r.hamza17):

Thanx alot I